Let alpha beta be the roots of ax^2 + bx + c is equal to zero and alpha1, - beta be the roots of a1x^2 + b1X + 7 is equal to zero then find a quadratic equation whose roots are  alpha and Alpha1?

Grade:10

1 Answers

Arun
5 years ago
Dear student
 
In first equation
 
alpha + beta = -b/a     (i)
alpha * beta = c/a       (ii)
 
Now
 
In second equation
 
alpha1 + ( - beta) = -b1/a1     (iii)
alpha1 *(-beta) = 7/a1             (iv)
 
Now add i and iii
 
alpha + alpha1 = -b/a +  -b1/a1
 
Also you can find product
 
Hope it helps